OpenStudy (anonymous):

What is the absolute value? |6-11|

OpenStudy (anonymous):

\[\left| 6-11 \right|=\left| -5 \right|=5\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

|3-2| ??

OpenStudy (anonymous):

absolute value is the distance from zero, so if 6-11 = -5, then it is 5 away from zero try the same thing with 3-2

OpenStudy (anonymous):

|3-2| = |1| = 1